Assessing the Impact of Suarez​ Manufacturing's Proposed Risky Investment on Its Stock Value   Early in​ 2013, Inez​ Marcus, the chief financial officer for Suarez​ Manufacturing, was given the task of assessing the impact of a proposed risky investment on the​ firm's stock value. To perform the necessary​ analysis, Inez gathered the following information on the​ firm's stock. During the immediate past 5 years​ (2008-2012), the annual dividends paid on the​ firm's common stock were as​ follows: Dividend per share 2012 $ 1.90 2011 1.70 2010 1.55 2009 1.40 2008 1.30 a. Find the current value per share of Suarez​ Manufacturing's common stock. The firm expects that without the proposed​ investment, the dividend in 2013 will be $ 2.09 per share and the historical annual rate of growth​ (rounded to the nearest whole​ percent) will continue in the future.​ Currently, the required return on the common stock is 14.0 % . ​Inez's research indicates that if the proposed investment is​ undertaken, the 2013 dividend will rise to $ 2.15 per share and the annual rate of dividend growth will increase to 13.0 % . She feels that in the best​ case, the dividend would continue to grow at this rate each year into the future and that in the worst​ case, the 13.0 % annual rate of growth in dividends would continue only through​ 2015, and​ then, at the beginning of​ 2016, would return to the rate that was experienced between 2008 and 2012. As a result of the increased risk associated with the proposed risky​ investment, the required return on the common stock is expected to increase by 2.0 % to an annual rate of 16.0 % ​, regardless of which dividend growth outcome occurs. Armed with the preceding​ information, Inez must now assess the impact of the proposed risky investment on the market value of​ Suarez's stock. To-Do A). Find the current value per share of Suarez​ Manufacturing's common stock.(B. Find the value of​ Suarez's common stock in the event that it undertakes the proposed risky investment and assuming that the dividend growth rate stays at 13.0 % forever. Compare this value to that found in part ​(a​). What effect would the proposed investment have on the​ firm's stockholders? Explain.C). On the basis of your findings in part ​(b​), do the stockholders win or lose as a result of undertaking the proposed risky​ investment? Should the firm do​ it? Why? D). Rework parts ​(a​) and ​(b​) assuming that at the beginning of 2016 the annual dividend growth rate returns to the rate experienced between 2008 and 2012. PLEASE ANSWER ALL PARTS TO THE QUESTIONS A THROUGH D

Please Anser all the questions this is my 3 rd time resubmitting!!!!!
Answer #1

The question gives data on the past and future dividends, the expected growth rates and the required return on equity. The method of valuing the stock that can be used is the Dividend Discount Model (DDM).

Lets go through the DDM in brief.

The value of any stock is the value of the total expected future cash flows. In this case the dividends that are given. In the question we also have a growth rate given for the dividends. So we need to calculate the dividends for each of the year. However, for the scenario where the dividend growth rate is constant forever, we can use the formula for valuation of perpetuity to arrive at the price of the stock. Pleas refer the formulas below.

D11 P = Tre -9

where,

P_{t} is Price per share of the stock at time t

g is the constant growth rate of the Dividends in perpetuity

D_{t+1} is the Dividend expected in time t+1 (i.e. expected in the next year) [D++1 = D+ * (1+9)

r_{e} is the required return on the stock/equity

Now lets look at each of the questions one by one.

A) Current value per share of Suarez​ Manufacturing's common stock

For this, we need to calculate the historical growth rate of the dividends as the company expects the dividends to grow at the same rate perpetually. This growth rate is the CAGR for the dividends between 2008 and 2012.

D_{2008} = $ 1.30, D_{2012} = $ 1.90, time t = 2012-2008 = 4 years

Therefore, constant growth rate, g = \left ( \frac{1.90}{1.30} \right )^{\tfrac{1}{4}} -1 = 10% (rounded to the next whole percent)

g = 10%, r_{e} = 14%, D_{2013} = $ 2.09 (details as given in the question for scenario with no investment proposal)

So, D2013 2.09 (0.14 - 0.10) 10 = 52.25 12 Tre-9)

So the current price per share of the stock is $ 52.25.

B) Proposed risky investment is undertaken and the growth rate stays at 13%

Under this scenario, the question states that the company expects the Dividend in 2013 to be $ 2.15

g = 13%, r_{e} = 16%, D_{2013} = $ 2.15 (details as given in the question for scenario with investment proposal)

So, D2013 2.15 (0.16 - 0.13) (re-g)

So the price per share of the stock with the risky investment proposal and the best case scenario i.e. growth rate at 13% perpetually is $ 71.67.

The proposed investment led to an increase stock price thereby benefiting the stockholders. This is attributed to two factors: a) Increase in the expected dividends for the next year i.e. 2013, b) Increase in the growth rate from 10% to 13% perpetually.

The major contributor here is the growth rate. This can be observed by calculating the stock price at an increase growth rate of 13% perpetually while still keeping the expected dividend for the year 2013 constant.

2.09 P2012 = 10.16 - 0.13) 0.16 - 0.13) = 69.67

It is can also be observed that the increase in the required return on the stock has a big negative impact on the stock price. Considering the dividend constant at $ 2.09, growth rate at 13% and a required return at 14% the stock price is

2.09 P2012 (0.14 - 0127 = 209 (0.14 -

C) On the basis of your findings in part ​(b​), do the stockholders win or lose as a result of undertaking the proposed risky​ investment? Should the firm do​ it? Why?

Looking at the results of the calculations done in part (B), the stockholders win as a result of the undertaking of the proposed risky investment. Considering only the scenario in part (B) and the results therein the firm should go ahead with the risky investment proposal. However, it is to be noted that this calculation is a reflection of the assumptions for a best case scenario.

The firm should decide on the risky investment post considering the worst case scenario and the probability of each of them occurring. The rework done in part (D) of the question with the assumptions for worst case scenario will help in understanding the impact of the proposed investment and give direction on the choice the firm should make.

D) Rework parts ​(a​) and ​(b​) assuming that at the beginning of 2016 the annual dividend growth rate returns to the rate experienced between 2008 and 2012.

Assumption for this part is that the dividend growth rate from 2016 onwards will return to the growth rate seen between 2008 and 2012. This has been calculated in part (A) to be 10%.

Please note that the calculations in part (A) were done with an assumption that the growth rate will be the same as the historical annual growth rate i.e. 10%. Therefore there is no rework needed on part(A) here. The price of the stock is $ 52.25.

For part (B), the growth rate returns to the rate between 2008 and 2012 from 2016. Therefore the dividends have two growth rates i.e. a high growth rate (g_{H}) phase from 2013 to 2015 and a constant growth rate (g_{c}) phase from 2016 perpetually.

Here, g_{H} = 13%, g_{c} = 10%, r_{e} = 16%

D_{2013} = $ 2.15 (given in question)

D_{2014} = $ 2.15 * (1+13%) = $ 2.43

D_{2015} = $ 2.43 * (1+13%) = $ 2.75

D_{2016} = $ 2.75 * (1+10%) = $ 3.03

D2013 D2014 D2015 P2015 I 2012 = (1+re)*(1+re)2 + (1+re)3* (1 + r.)3

D2016 D2013 (1+r) D2014 (1 + r.)2 - D2015 (1+re)3 + (-9) (1 + r.)3 2012

3.03 2.15 2.43 2.75,0.16-0.10) -37.77 2012 = 116 + 1.162 + 1.168 + 1.163

So, in the worst case scenario the risky investment reduces the value of the stock.

P_{2012}(no investment) = $ 52.25

P_{2012}(with investment and best scenario) = $ 71.67

P_{2012}(with investment and worst scenario) = $ 37.77

From the above prices we see that the proposed risky investment has a potential upside of +37.17% and a potential downside of -27.71% impact on the current stock price. Therefore the probability of each of the scenarios need to be considered before taking a decision on the risky investment.